I have four questions I'm experiencing difficulty with:
1) The difference of the product of 3 and a number and -1/2 is -5/2.
2) three less than the quotient of a number and 2 is 8.
3) The product of a number divided by 5 and 2 is 1.
4) The sume of 2 and 3, plus 3 times a number, is -4.
5) the quotient of -36 and 9, subtracted from 4 times a number, is -8.
Thank you in advance for your help.

Answer by edjones(8007)

1) let the number be x.
3x-(-1/2)=-5/2
3x+1/2=-5/2
3x+1/2-1/2=-5/2-1/2
3x=-6/2
3x=-3
x-1
.
2)x/2 -3=8
2(x/2 -3)=2*8
x-6=16
x-6+6=16+6
x=22
.
3)2x/5=1
5*2x/5=1*5
2x=5
2x/2=5/2
x=5/2
.
4) 2+3+3x=-4
3x+5=-4
3x+5-5=-4-5
3x=-9
3x/3=-9/3
x=-3
.
I'll leave #5 up to you.
.
Ed